OpenStudy (loser66):

Let R be a ring and X is a set. a) Show that the set Fun(X,R) of functions on X with values in R is a ring. b) Show that R is isomorphic to the subring of constant functions on X. c) Show that Fun(X, R) is commutative iff R is commutative. d) Suppose that R has an identity, show that Fun (X,R) has an identity and describe the units of Fun(X,R) Please, help

OpenStudy (loser66):

@zzr0ck3r

OpenStudy (zzr0ck3r):

start with a?

OpenStudy (loser66):

If you can, please give me all. I missed only 1 class and ..... blank

OpenStudy (zzr0ck3r):

I am going to look at it on paper, and what ever I figure out I will post the paper.

OpenStudy (loser66):

Thanks a ton.

OpenStudy (zzr0ck3r):

But give me a few, I am in topology mode.... so it could take a second to make the switch lol.

OpenStudy (loser66):

I am not in hurry, whenever you have time. :) (but not after Monday, I have another thing on Monday, can't miss this part before getting a new one, please)

OpenStudy (zzr0ck3r):

so do you know the operations on the functions? composition and addition?

OpenStudy (loser66):

No :(

OpenStudy (zzr0ck3r):

I am guessing it is...

OpenStudy (zzr0ck3r):

hmm

OpenStudy (loser66):

Can you give me any link to read for this part? I don't want to be blink like this. I want to know .....all. :)

OpenStudy (zzr0ck3r):

what book are you using?

OpenStudy (loser66):

The free book online Algebra abstract and concrete 2.6edition http://homepage.math.uiowa.edu/~goodman/22m121.dir/2005/sections6.1-6.3.pdf part 6.1, 6.2 but I don't see the notation Fun (X,R) on it.

OpenStudy (loser66):

and the problem is 6.1.12

OpenStudy (zzr0ck3r):

a) For a ring \((R,+,*)\) we will use the inherited operations for the functions. i.e. \((f+g)(x)=f(x) +g(x)\text{ and } (f*g)(x) = f(x)*g(x)\). Let \(X\subset R\). Consider the set \(\text F=\text{Fun}(X,R)\). Let \(f\in F \), and consider \(f_0:X\rightarrow R, f(x)=0\) where \(0\) is the additive identity in \(R\). Then \((f+f_0)(x) = f(x)+f_0(x) =f(x) + 0 =f(x) \ \ \ \forall \ x\in X \). Similarly \((f_0+f)(x) = f(x)\). So we have additive identity in \(F\). if \(f:X\rightarrow R,f(x)=a\) is a in \(F\), then so is \(g:X\rightarrow R,g(x)=-a\) is also in \(F\). So \((f+g)(x) = a-a=0 \ \ \ \ \forall x\in X\). So \(g=-f\) So we have inverses. Let \(f, g, h \in F\) then since there image is in \(R\), \((f+g)+h = f+(g+h)\), Similarly \(f+g=g+f\). So \(F\) forms a group and we are halfway done with a). you with me?

OpenStudy (zzr0ck3r):

Can you show that 1) \(F\) is associative under multiplication (given that R is). 2) there is an identity function (given \(1\in R\)). note, it's not f(x) = x. 3) \(f*(g+h) = f*g+f*h\) for all \(x\). Then we are done with part a). Since we are not in a field we don't need multiplicative inverses...

OpenStudy (zzr0ck3r):

I know you are having inet probs, so I will wait before I go on. Since you are in no rush and I will be around for the next few days ;)

OpenStudy (loser66):

For your questions: 1) Let f, g, h ∈F , we have f(g(h(x) ) = fg(h(x) = fgh(x) , the same for approaching fro the right. 2) To be a ring, set doesn't need to have multiplicative identity, right? but if needed, the multiplication identity of F should be \(f_i =1\) 3) honestly, to me, it's the hardest part when I have to prove the trivial problem. Is it not that (g+h)(x) = g(x) +h(x) and f(g+h) (x) = f*g+f*h

OpenStudy (zzr0ck3r):

Some definitions include multiplicative identity and some don't...So its fine to skip that part if you are working in a book that does not. I actually think they wanted you to use the multiplication inherited from R, check out how I defined multiplication on \(F\). In this case, since multiplication distributes in \(R\), function multiplication is trivial(the codomain is R, so your image is in R).

OpenStudy (zzr0ck3r):

For the isomorphism Let \(C\) be the set of constant functions, then \(g:R\rightarrow C, g(a) = f_a\) where \(f_a\) is the constant function with output \(a\), is an isomorphism.

OpenStudy (zzr0ck3r):

For commutative property, note that \(\forall \ f \in F \ \forall \ x\in X \ (f(x) \in R )\ \), so given any \(f,g\in F\) we have \( \ (f*g)(x) = f(x)*g(x)=g(x)*f(x)=(g*f)(x) \ \forall \ x\in X\).

OpenStudy (zzr0ck3r):

So \(R\) being commutative implies \(F\) is commutative. Suppose \(F\) is commutative, than any subring is commutative. We proved \(C\) was a subring by showing it was isomorphic to ring \(R\)(surely its a subset of \(F\)). So \(R\) is commutative because \(C\) is a commutative subring.

OpenStudy (zzr0ck3r):

Its easy to show that \(f_1\in C\) is the identity for multiplication in \(F\).

OpenStudy (zzr0ck3r):

Up there where it says In this case, since multiplication distributes in \(R\) , function multiplication is trivial(the codomain is \(R\), so your image is in \(R\)). it should say In this case, since multiplication distributes in R , function multiplication distributes(the codomain is R, so your image is in R).

OpenStudy (zzr0ck3r):

how we doing?

OpenStudy (loser66):

for the last part, suppose F has identity, by definition of the unit in ring, we have set of the units in F(X,R) = { g(x) | g(x) = \(f^{-1}(x)\) x \(\in X\)}, right?

OpenStudy (zzr0ck3r):

yeah that's fine

OpenStudy (zzr0ck3r):

Make sure that you note that you mean multiplicative inverse not function inverse i.e. \(f^{-1}=\frac{1}{f}\)

OpenStudy (loser66):

Question on part b) C is a subring of R, hence g is one of endomorphism from R to itself, right? all we need to prove is g is bijective to get g is isomorphism. Am I right?

OpenStudy (loser66):

for d) show that F has an identity Question: from a) we proved that F is a ring, so that F is an addition abelian GROUP, which always has identity. All we need is to define multiplication identity, right?